Please help me solve this: the three angles in a triangle are in the extended ratio 3:9:12, Find each.
Answer by TimothyLamb(4379)

x, y, and z = the measures of the three angle, in degrees:
---
3/9 = x/y
9/12 = y/z
x + y + z = 180
---
put the system of linear equations into standard form
---
9x - 3y = 0
12y - 9z = 0
x + y + z = 180
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 22.5
y = 67.5
z = 90
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---